# Mobile-Efficient Speech Emotion Recognition Using DistilHuBERT: A Cross-Corpus Validation Study

### v1

@@ VI Conclusion
Cross-corpus evaluation on RAVDESS further revealed the model’s strengths and limitations. While overall accuracy dropped to 43.29% due to domain shift, the error analysis demonstrated that misclassifications were not random but systematically clustered by arousal level. The model maintained 97% recall for anger and 64% recall for sadness, indicating robust detection of physiological arousal states despite the acoustic mismatch between semi-spontaneous and theatrical speech. This behavior suggests that the model has learned generalizable arousal-related features, even if valence discrimination remains sensitive to recording environment and expressive style.
@@
Future work should explore several directions to further improve mobile SER systems. First, the theatricality gap identified in our cross-corpus analysis suggests that domain adaptation techniques, such as adversarial training or style transfer, could help align representations across corpora with differing expressive intensities. Second, the proposed VAD-based deployment pipeline for long-form voice notes requires empirical validation to quantify its effectiveness in reducing neutral-class bias. Third, extending the model to support fine-grained dimensional emotion prediction (arousal and valence) rather than categorical labels could provide more nuanced affect recognition while leveraging the model’s demonstrated strength in arousal detection. Finally, investigation of model pruning and neural architecture search may reveal even more compact architectures without further sacrificing accuracy.
In conclusion, this work establishes that distilled transformer models, when combined with rigorous evaluation protocols and strategic cross-corpus training, can bridge the gap between laboratory performance and practical mobile deployment. The resulting system provides a foundation for privacy-preserving, low-latency emotion recognition in real-world applications, from mental health monitoring to adaptive conversational agents.
@@ Acknowledgment
All experimental code and trained models are publicly available to facilitate reproducibility and future research.
Appendix A Reproducibility: Experimental Notebooks
@@
TABLE III: Publicly Available Experimental Notebooks
@@
5-fold LOSO CV on IEMOCAP. Baseline in Table I.
kaggle.com/code/amrproject5/distilhubert-on-iemocap
@@
First 3 folds with CREMA-D. See Table I.
kaggle.com/code/amrproject5/distilhubert-on-crema-d-and-iemocap
@@
Final 2 folds with CREMA-D. See Table I.
kaggle.com/code/saifeldenmohamed8/distilhubert-on-iemocap-and-crema-d-2
@@
kaggle.com/code/saifeldenmohamed8/production-script
@@
kaggle.com/code/saifeldenismail/cross-corpus-analysis

### v2

@@ VI Conclusion
Cross-corpus evaluation on RAVDESS further revealed the model’s strengths and limitations. While overall accuracy dropped to 46.64% due to domain shift, the error analysis demonstrated that predictions clustered by arousal level rather than by specific emotion categories—happiness bled into anger, and sadness bled into neutral. The model maintained 99% recall for anger, 55% recall for neutral, and 27% recall for sadness, indicating robust detection of physiological arousal states despite the acoustic mismatch between semi-spontaneous and theatrical speech. This behavior suggests that the model has learned generalizable arousal-related features, even if valence discrimination remains sensitive to recording environment and expressive style.
@@
Future work should explore several directions to further improve mobile SER systems. First, the theatricality gap identified in our cross-corpus analysis suggests that domain adaptation techniques, such as adversarial training or style transfer, could help align representations across corpora with differing expressive intensities. Second, the proposed Voice Activity Detection-based deployment pipeline for long-form voice notes requires empirical validation to quantify its effectiveness in reducing neutral-class bias. Third, extending the model to support fine-grained dimensional emotion prediction (arousal and valence) rather than categorical labels could provide more nuanced affect recognition while leveraging the model’s demonstrated strength in arousal detection. Finally, investigation of model pruning and neural architecture search may reveal even more compact architectures without further sacrificing accuracy.
In conclusion, this work demonstrates that distilled transformer models, when combined with rigorous evaluation protocols and strategic cross-corpus training, can bridge the gap between laboratory performance and practical mobile deployment. The resulting system provides a foundation for privacy-preserving, low-latency emotion recognition in real-world applications, from mental health monitoring to adaptive conversational agents.
@@ Acknowledgment
All experimental code and trained models are publicly available to facilitate reproducibility and future research. This work received no specific funding from agencies in the public, commercial, or not-for-profit sectors.
Appendix A Reproducibility: Experimental Notebooks and Links to Code
@@
TABLE III: Publicly Available Experimental Notebooks and Links to Code
@@
5-fold LOSO CV on IEMOCAP. Baseline in Table II.
https://www.kaggle.com/code/amrproject5/distilhubert-on-iemocap
@@
First 3 folds with CREMA-D. See Table II.
https://www.kaggle.com/code/amrproject5/distilhubert-on-crema-d-and-iemocap
@@
Final 2 folds with CREMA-D. See Table II.
https://www.kaggle.com/code/saifeldenmohamed8/distilhubert-on-iemocap-and-crema-d-2
@@
https://www.kaggle.com/code/saifeldenmohamed8/production-script
@@
https://www.kaggle.com/code/saifeldenismail/cross-corpus-analysis